# Language Model Agents Suffer from Compositional Generalization in Web Automation

### v1

@@ 4.1 RCI
4.2 AdaPlanner
In contrast to other prompted LMAs, AdaPlanner (Sun et al., 2023) leverages the capability of program synthesis in LLMs to mitigate the hallucination in a plan. Conditioning on the instruction, the description of permissible actions in the web environments, and few-shot demonstrations, the agent first generates an open-loop plan in a Python function, where each snippet corresponds to the action. Once the agent receives environmental feedback at each step, such as assertion errors in the code, other functional errors, or “ask” action to LLMs, it adaptively re-generates the plan for the remaining steps in a closed-loop manner. It has been reported that LLMs more capable of code generation perform better, such as text-davinci-003 than gpt-3.5-turbo.
@@ 4.3 Synapse
Synapse (Zheng et al., 2023) argues that LMAs perform better if well-designed structured prompts are provided, even without self-improvement or program synthesis. The structured prompting is formed by two pre-processing strategies: state filtering and task reformulation. State filtering gradually transforms raw HTML into simple formatted text, such as a Pythonic list or dictionary, in a multi-step manner, which may improve the state understanding of LMAs. Task reformulation translates given instructions or raw HTML into decomposed queries: for instance, translating “select 12/03/2016 as the date and hit submit” into “select the datepicker at step 1, click ’Prev’ 7 times at step 2-8 (May is 7 months before December), click the date ’12’ at step 9, and finally submit at step 10” (translated instruction), or mapping proper noun into corresponding XPath (translated HTML).
While detailed structured prompts have led to strong performances, those should be specialized for each primitive task in MiniWoB by leveraging 7 different types of reformulation. See Appendix C.3 for further details.
@@ 4.5 Data-Rebalancing Improves Finetuned Language Model Agents
Furuta et al. (2023) utilized agent-driven data collection instead of humans to improve the performance of finetuned LMAs further. For each task, 10k demonstrations are collected and filtered based on task success, which resulted in challenging tasks having much less than 10k demonstrations due to the sub-optimal performance of LMA on these tasks. We identify that by fixing the data-imbalance problem, the performance of finetuned LMAs can be significantly improved, achieving super-human performance on MiniWoB. We first run Synapse (Zheng et al., 2023) on MiniWoB and collect 77K additional demonstrations across 16 tasks on top of 347K demonstrations (Furuta et al., 2023) to compensate for the lack of data in specific tasks. We then estimate the “brute-force” task difficulty averaging success rates for representative web automation agents. Based on those proximal measures, we classify 65 base tasks into three categories, such as easy (0.8 - 1.0), medium (0.6 - 0.8), and hard (0.0 - 0.6) (see Appendix D). We then balance the number of episodes based on the task difficulty, where we gradually reduce the ratio of easier tasks to focus more on challenging tasks. For instance, we remove $X$episodes from top-$k$ tasks in easy group (see Appendix E for the details).
@@
We finetune HTML-T5-XL (Gur et al., 2023), a pre-trained language model with local and global attention in the encoder and a mixture of long-span denoising, on these rebalanced datasets. Table 1 shows that all the data-rebalance strategies improve the success rate, and reducing 50% episodes from easy tasks (finally 282K episodes in total) is the most effective rebalancing strategy. This suggests that finetuned LMAs can be as capable as prompted LMAs in decision making tasks. We include HTML-T5++ as a baseline in the following sections.
5 Designing CompWoB to Measure Compositional Generalization
@@
Even though MiniWoB includes a spectrum of simulated environments, they have still focused on narrow and single-task instances. We need more advanced environments to measure the generalization to the various challenges in real-world web automation, such as complex observation (Deng et al., 2023), instruction (Zhou et al., 2023b), and task compositionality (Gur et al., 2021). We design CompWoB (i.e. compositional MiniWoB), as a novel test bed for LMAs, by leveraging the high composability of simulated web environments (Figure 1). In CompWoB, we systematically combine several base tasks (from 2 to 8) in the original MiniWoB, which LMAs can already solve, into a single task (e.g. click-checkboxes-transfer + enter-password + click-dialog $\rightarrow$ click-checkboxes-transfer_enter-password_click-dialog). This allows us to control the complexity of HTML and instructions, ensuring novel tasks are solvable to some extent. CompWoB includes realistic task compositions, such as combining form filling, popup message, and login page, in a single or across multiple pages.
As mentioned in Section 4.5, we first calculate the average success rates among representative web automation agents as brute-force task complexity, and classify 65 primitive tasks in MiniWoB into 3 categories (easy, medium, hard) based on those scores. The details of classification are described in Appendix D. We randomly select base tasks from easy group, filter those combinations by their feasibility, and make 50 compositional tasks. We divide those into five categories: two-way tasks (20), three-way tasks (10), n-way tasks (5), transition tasks (5), and easy-medium two-way tasks (10). In n-way tasks, we combine from 4 to 8 tasks sequentially, and in transition tasks, we implement explicit page transition, for instance, transiting from the login form to the email browser. We also sample several tasks from medium group to construct easy-medium two-way tasks. You can find the full list of tasks in Appendix H. We simply stitch the instructions with “and then” and put each HTML on the same depth. LMAs should satisfy the given instructions sequentially, such as from task A, B, to C. Moreover, to test whether LMAs can deal with complex and ambiguous instructions, we propose reverse-order instruction settings, where the instruction is provided upside down while its task order is semantically the same (e.g. solve task B and C, after solving A). These simple yet controllable strategies make the analysis of LMA’s behaviors tractable while reflecting compositional aspects of real-world tasks.
@@ 6 Results
Evaluation Methodology  We evaluate both transferred and prompted LMAs with base MiniWoB demonstrations on the unseen compositional tasks in a “zero-shot” manner; i.e. we do not provide any demonstrations on the compositional tasks for the training corpus and exemplars to measure the generalization. We test 50 compositional tasks and run 100 episodes per task. We adopt gpt-3.5-turbo as a backbone LLM, unless otherwise mentioned. We assume the optimal exemplar retriever throughout experiments and always provide the pre-defined prompts to LMAs. We borrow hyper-parameters and prompt templates from respective papers with minimal change to respect our zero-shot transfer setting.
@@
AdaPlanner  Following the original implementation, we use the exemplars provided by Sun et al. (2023) for tasks where those base tasks are included, such as enter-text and click-widget (see Appendix C.2). Otherwise, the agents are prompted in a zero-shot manner.
Synapse  We test 3 prompting strategies: few-shot with (1) first-task exemplars, (2) second-task exemplars, and (3) best exemplars (i.e. maximum score between (1) and (2)). Because prompts and modules are quite different among the primitive tasks, we do not merge the prompts and just use proper hyper-parameters corresponding to the given exemplars designed by Zheng et al. (2023).
Figure 3: Average success rate of language model agents in reverse-order instruction settings. We use gpt-3.5-turbo as the backbone LLM for prompted LMAs (RCI, AdaPlanner, Synapse), and transferred LMAs with 3-billion parameters. Notably, most LMAs significantly degrade the success rate when reverse-order instructions are provided. This trend is more remarkable in transferred LMA (from 54.8% to 31.0% on average) than prompted LMA (from 24.9% to 18.0% on average), which suggests that any kind of LMAs are susceptible to the order of compositional instructions. The capability as general language models might be important to parse semantically complex instructions into the correct plan.
6.1 Language Model Agents Struggle to Handle Task Compositionality
Figure 2 shows that, in CompWoB, all the LMAs face performance degradation. Among those, transferred LMAs achieve better success rate (54.8%) than prompted LMAs (24.9%) on average. In particular, HTML-T5++ achieves the best generalization while suppressing the performance drop from 95.2% to 61.5%. In contrast, prompted LMAs degrade their performance drastically; even the best RCI with few-shot combination exemplars (comb) just degrades the success rate to 28.7% from 90.6% in base MiniWoB. These results indicate that LMAs suffer from generalization to task compositionality, and transferred LMAs can relatively deal with that better than prompted LMAs, which is an opposite trend to base MiniWoB performance. Among prompted LMAs, RCI performs better than AdaPlanner and Synapse, which suggests that multiple iterations of self-criticism and improvement might be more robust to out-of-domain decision making from the exemplars than program synthesis with feedback or structured prompting with state translations.
In the failure episodes (Table 2), LMAs often miss necessary steps, common to all the prompted LMAs. Since the instructions get long in compositional settings, LMAs may skip important intermediate steps to satisfy the instructions. In addition, they predict incorrect action types and XPath: for instance, hallucination in XPath (RCI) and mixing up click and type action (Synapse).
@@ 6.1 Language Model Agents Struggle to Handle Task Compositionality
Table 2: Failure examples in CompWoB. The left columns have correct plans and the right columns have failure plans. LMAs often ignore necessary intermediate steps or predict incorrect action types and XPath.
@@ 6.2 Reverse-Order Instructions Degrade Language Model Agents
As shown in Figure 3, all the LMAs significantly degrade the success rate when reverse-order instructions are provided. This trend is more remarkable in transferred LMAs dropping from 54.8% to 31.0% than prompted LMAs from 24.9% to 18.0% on average, which suggests that any kind of LMAs is susceptible to the order of compositional instructions and that transferred LMAs may not generalize well to diverse instructions beyond the dataset distribution. As opposed to Section 6.1, the performance differences among prompted LMAs are marginal, which implies that existing prompting methods, even with self-improvement, may not handle complex task instructions enough. The stronger capability as general-purpose language models or other prompting methods might be important to parse semantically complex instructions into the executable sequential plan.
Compared to Section 6.1, LMAs cannot parse reverse-order instructions into plans correctly (Table 3), which is observed with RCI and Synapse but not with AdaPlanner. LMAs still fail to select correct action types (AdaPlanner) or XPath (Synapse), and they also predict unnecessary actions (RCI).
@@
Figure 4: Average success rate of large language models with advanced LLMs (gpt-4 for RCI and Synapse, and text-davinci-003 for AdaPlanner). The lighter color represents the performance in MiniWoB, the medium color does in CompWoB with gpt-3.5-turbo, and the darker color does with gpt-4 or text-davinci-003. The more capable models (gpt-4 as a generalist and text-davinci-003 as a coder) can improve the success rate of prompted LMAs but still struggle to generalize to compositional tasks (e.g. 56.0% by RCI) or to deal with reverse-order instructions (e.g. 43.5% by RCI). This may indicate that we need much better foundation models to realize deployable LMA in the complex real world.

### v3

@@ 4.3 Synapse
Synapse (Zheng et al., 2023) argues that LMAs perform better if well-designed structured prompts are provided, even without self-improvement or program synthesis. The structured prompting is formed by two pre-processing strategies: state filtering and task reformulation. State filtering gradually transforms raw HTML into simple formatted text, such as a Pythonic list or dictionary, in a multi-step manner, which may improve the state understanding of LMAs. Task reformulation translates given instructions or raw HTML into decomposed queries: for instance, translating “select 12/03/2016 as the date and hit submit” into “select the datepicker at step 1, click ’Prev’ 7 times at step 2-8 (May is 7 months before December), click the date ’12’ at step 9, and finally submit at step 10” (translated instruction), or mapping proper noun into corresponding XPath (translated HTML). While detailed structured prompts have led to strong performances, those should be specialized for each primitive task in MiniWoB by leveraging 7 different types of reformulation. See Appendix C.3 for further details.
@@ 4.5 Data-Rebalancing Improves Finetuned Language Model Agents
Furuta et al. (2023) utilized agent-driven data collection instead of humans to improve the performance of finetuned LMAs further. For each task, 10k demonstrations are collected and filtered based on task success, which resulted in challenging tasks having much less than 10k demonstrations due to the sub-optimal performance of LMA on these tasks. We identify that by fixing the data-imbalance problem, the performance of finetuned LMAs can be significantly improved, achieving super-human performance on MiniWoB. We first run Synapse (Zheng et al., 2023) on MiniWoB and collect 77K additional demonstrations across 16 tasks on top of 347K demonstrations (Furuta et al., 2023) to compensate for the lack of data in specific tasks. We then estimate the “brute-force” task difficulty averaging success rates for representative web automation agents. Based on those proximal measures, we classify 65 base tasks into three categories, such as easy (0.8 - 1.0), medium (0.6 - 0.8), and hard (0.0 - 0.6) (see Appendix D). We then balance the number of episodes based on the task difficulty, where we gradually reduce the ratio of easier tasks to focus more on challenging tasks. For instance, we remove $X$episodes from top-$k$ tasks in easy group. We heuristically design the following data-mixing strategies:
  • Original dataset released by Furuta et al. (2023) (347K episodes).
  • Adding 77K episodes from 16 tasks suffering the lack of data (424K episodes).
  • Removing 50% episodes from top-10 easy tasks (424K - 73K = 351K episodes)
  • Removing 80% episodes from top-10 easy tasks (424K - 102K = 322K episodes)
  • Removing 50% episodes from easy tasks (424K -142K = 282K episodes)
  • Removing 80% episodes from top-15 easy tasks and removing 50% episodes from other 11 easy tasks (424K - 183K = 241K episodes)
See Appendix E for further details. The automation of data mixture design (Xie et al., 2023) would be an important future direction.
@@
We finetune HTML-T5-XL (Gur et al., 2023), a pre-trained language model with local and global attention in the encoder and a mixture of long-span denoising, on these rebalanced datasets. Table 1 shows that all the data-rebalance strategies improve the success rate, and reducing 50% episodes from easy tasks (finally 282K episodes in total) is the most effective rebalancing strategy. This suggests that finetuned LMAs can be as capable as prompted LMAs in decision making tasks. Moreover, due to the inherent compositional nature in challenging web automation tasks, data rebalancing may improve the capability of dealing with unseen task compositions. We include HTML-T5++ as a baseline in the later sections.
# of Episode Steps # of Instruction Tokens
Benchmark Average 50th 90th 95th Max Average 50th 90th 95th Max
MiniWoB (Shi et al., 2017) 3.6 4.3 5.5 5.8 7.0 8.5 14.8 30.5 12.9 23.2
Mind2Web (Deng et al., 2023) 7.7 6.0 14.0 17.0 37.0 19.2 16.0 33.0 42.0 84.0
CompWoB (Ours) 7.2 6.8 12.8 14.2 17.8 37.3 34.2 55.2 57.9 100.0
Table 2: The statistics (Mean, Max, 50/90/95th percentiles) of episode steps and instruction tokens from MiniWoB (Shi et al., 2017), Mind2Web (Deng et al., 2023), and CompWoB. CompWoB successfully increases task complexity from MiniWoB, and has sufficient difficulties the same as Mind2Web, a representative benchmark from real websites.
5 Designing CompWoB Controlling Task Complexity and Compositionality
Even though MiniWoB includes a spectrum of simulated environments, they have still focused on narrow and single-task instances. We need more advanced environments to measure the generalization to the various functional challenges in real-world web automation, such as complex observation (Deng et al., 2023), instruction (Zhou et al., 2023b), and task compositionality (Gur et al., 2021). We design CompWoB (i.e. compositional MiniWoB), as a general test bed for LMAs, by leveraging the high composability of simulated web environments (Figure 1). In CompWoB, we systematically combine several base tasks (from 2 to 8) in the original MiniWoB, which LMAs can already solve, into a single task (e.g. click-checkboxes-transfer + enter-password + click-dialog $\rightarrow$ click-checkboxes-transfer_enter-password_click-dialog). This allows us to control the complexity of HTML and instructions, ensuring novel tasks are solvable to some extent. Moreover, CompWoB can work as a hold-out test environment accompanying with MiniWoB as a train environment. While tasks are visually simplified, CompWoB reflects a broader spectrum of functionalities in real-world websites without sacrificing controllability. For instance, the task in Figure 1 sketches the structure of a login form with agreement checkboxes and a pop-up window, like Internet banking. We also designed tasks with page transitions, such as from the login form to the email browser.
5.1 Details of Task Design
As mentioned in Section 4.5, we first calculate the average success rates among representative web automation agents as brute-force task complexity, and classify 65 primitive tasks in MiniWoB into 3 categories (easy, medium, hard) based on those scores. The details of classification are described in Appendix D. We randomly select base tasks from easy group, filter those combinations by their feasibility and reality, and make 50 compositional tasks. We divide those into five categories: two-way tasks (20), three-way tasks (10), n-way tasks (5), transition tasks (5), and easy-medium two-way tasks (10). In n-way tasks, we combine from 4 to 8 tasks sequentially, and in transition tasks, we implement explicit page transition, for instance, transiting from the login form to the email browser. We also sample several tasks from medium group to construct easy-medium two-way tasks. You can find the full list of tasks in Appendix H. For the implementation, we stitch the instructions with “and then” and put each HTML on the same depth. LMAs should satisfy the given instructions sequentially, such as from task A, B, to C. Moreover, to test whether LMAs can deal with complex and ambiguous instructions, we propose reverse-order instruction settings, where the instruction is provided upside down while its task order is semantically the same (e.g. solve task B and C, after solving A). These simple yet controllable strategies make the analysis of LMA’s behaviors tractable while reflecting compositional aspects of real-world tasks.
5.2 Number of Tasks
MiniWoB has around 104 base tasks in total; all the two-way and three-way combinations of these tasks would give 185,460 compositional tasks. It is infeasible to manually curate all the two-way/three-way combinations. Unfortunately, it is also nontrivial to automate this process due to the locality of the environment implementations in MiniWoB. Each environment is mostly self-contained, making it nontrivial to modularize the whole benchmark/codebase so that every combination can be automatically generated. Alternatively, we outline a set of design principles – solvability, feasibility, and reality, which we follow to manually curate 50 compositional tasks and 50 reverse-order instruction tasks that could cover a wide variety of difficulty, and compositionality. We believe these guiding principles would be applicable to other compositional generalization problems such as robotic navigation. Based on these design principles, our compositional benchmark can easily be extended in the future to study even more challenging and compositional web automation problems. We would also like to highlight that, because previous works were tested around 50 - 60 MiniWoB tasks (Kim et al., 2023; Sun et al., 2023; Zheng et al., 2023; Gur et al., 2022; Furuta et al., 2023), 50 compositional tasks is a decent number of tasks to evaluate the agents. Our addition of a hold-out compositional test set doubles the number of tasks for the community to study.
5.3 Inherent Compositionality and Sub-Task Dependency in Web Automation
While many of the real-world tasks have inherent compositionality to some degree, these tasks are not explicitly designed for compositionality, making it difficult to systematically investigate the generalization gap. This can be analyzed with our CompWoB by separating the difficulty of the base tasks themselves and the difficulty of the task compositions.
Real-world websites have dependencies between sub-tasks. For instance, an email client requires a successful login to proceed to writing an email or reading social media posts of a specific user requires navigating to the personal page of that user. In CompWoB, we have implemented these kinds of dependencies in the reward function using logical AND operations. For instance, while we allowed agents to continue to “write an email” sub-task without successful login, the episode is flagged as a failure (i.e. zero reward) even if the “write an email” sub-task is successful. We inform agents of these dependencies using connectors in instructions such as “solve B after A” or “solve A then B”. Our analysis in Section 6.4 also suggests that, in addition to task compositionality, long instructions and deep HTML sub-tree can lead to challenging tasks.
5.4 Comparison to Other Web Automation Benchmarks
We design CompWoB on top of MiniWoB, which enables us to test the human-level LMAs (Kim et al., 2023; Sun et al., 2023; Zheng et al., 2023; Gur et al., 2023) as strong baselines. Because the success of in-domain tasks is ensured, we could focus on the analysis of generalization to unknown task compositions and their functionality in a controllable way. In contrast, other benchmarks copying real websites fail to obtain decent agents (for instance, the episode success rate of Mind2Web (Deng et al., 2023) is around 5-10% at most; around 15% success at most in WebArena (Zhou et al., 2023b)) and it seems to be harder to identify the attribution of errors due to the complex nature of real websites. Moreover, while realistic benchmarks (Deng et al., 2023; Zhou et al., 2023b) often provide deterministic instructions and observations, CompWoB has randomized instructions generated from a set of templates, as well as element-randomized HTML. For the task complexity compared to other benchmarks, we provide the statistics (mean, max, percentiles) of episode steps and instruction tokens from MiniWoB (Shi et al., 2017), Mind2Web, and CompWoB in Table 2. In terms of episode length and instruction length, CompWoB successfully increases task complexity from MiniWoB, and has sufficient difficulties the same as Mind2Web, a representative benchmark from real websites.
@@ 6 Results
Evaluation Methodology  We evaluate both transferred and prompted LMAs with base MiniWoB demonstrations on the unseen compositional tasks in a “zero-shot” manner; i.e. we do not provide any demonstrations on the compositional tasks for the training corpus and exemplars to measure the generalization. We test 50 compositional tasks and run 100 episodes per task. We adopt gpt-3.5-turbo as a backbone LLM, unless otherwise mentioned. We assume the optimal exemplar retriever throughout experiments and always provide the pre-defined prompts to LMAs. We borrow hyper-parameters and prompt templates from respective papers with minimal change to respect our zero-shot transfer setting. As a sanity check for the quality of environments and baseline agents, we also test LMAs with oracle demonstrations in Appendix I.
@@
AdaPlanner  Following the original code, we use the same exemplars provided by Sun et al. (2023) for tasks where those base tasks are included, such as enter-text and click-widget (see Appendix C.2). Otherwise, the agents are prompted in a zero-shot manner.
Synapse  We test 3 prompting strategies: few-shot with (1) first-task, (2) second-task exemplars, and (3) best exemplars (i.e. maximum score between (1) and (2)). Because prompts and modules are quite different among tasks, we do not merge the prompts and use proper hyper-parameters corresponding to the given exemplars designed by Zheng et al. (2023).
@@
Table 3: Failure examples in CompWoB. The left columns have correct plans and the right have failure plans. LMAs often ignore necessary intermediate steps or predict incorrect action types and XPath.
6.1 Language Model Agents Struggle to Handle Task Compositionality
Figure 2 shows that, in CompWoB, all the LMAs face performance degradation. Among those, transferred LMAs achieve better success rate (54.8%) than prompted LMAs (24.9%) on average. In particular, HTML-T5++ achieves the best generalization, suppressing the performance drop from 95.2% to 61.5%. This might be because finetuned LLMs are better at natural language compositional tasks prompted LLMs in general. See Appendix K for further details. In contrast, prompted LMAs degrade their performance drastically; even the best RCI with few-shot combination exemplars (comb) just degrades the success rate to 28.7% from 90.6% in base MiniWoB. These results indicate that LMAs suffer from generalization to task compositionality, and transferred LMAs can relatively deal with that better than prompted LMAs, which is an opposite trend to base MiniWoB performance. Among prompted LMAs, RCI performs better than AdaPlanner and Synapse, which suggests that multiple iterations of self-criticism and improvement might be more robust to out-of-domain decision making from the exemplars than program synthesis with feedback or structured prompting with state translations.
In the failure episodes (Table 3), LMAs often miss necessary steps, common to all the prompted LMAs. Since the instructions get long in compositional settings, LMAs may skip important intermediate steps to satisfy the instructions. In addition, they predict incorrect action types and XPath: for instance, hallucination in XPath (RCI) and mixing up click and type action (Synapse). Without oracle exemplars, LMAs often struggle to parse compositional task instructions and understand complex HTML. Appendix H also provides average success rates per category.
Figure 3: Average success rate of language model agents in reverse-order instruction settings. We use gpt-3.5-turbo as the backbone LLM for prompted LMAs (RCI, AdaPlanner, Synapse), and transferred LMAs with 3-billion parameters. Notably, most LMAs significantly degrade the success rate when reverse-order instructions are provided (the performance gap is highlighted with a red number). This trend is more remarkable in transferred LMA (54.8% $\rightarrow$ 31.0% on average) than prompted LMA (24.9% $\rightarrow$ 18.0% on average), which suggests that any kind of LMAs are susceptible to the order of compositional instructions. The capability as general language models might be important to parse semantically complex instructions into the correct plan.
@@ 6.2 Reverse-Order Instructions Degrade Language Model Agents
As shown in Figure 3, all the LMAs significantly degrade the success rate when reverse-order instructions are provided. This trend is more remarkable in transferred LMAs dropping from 54.8% to 31.0% than prompted LMAs from 24.9% to 18.0% on average, which suggests that any kind of LMAs is susceptible to the order of compositional instructions and that transferred LMAs may not generalize well to diverse instructions beyond the dataset distribution. As opposed to Section 6.1, the performance differences among prompted LMAs are marginal, which, however, implies that existing prompting methods, even with self-improvement, may not handle complex task instructions enough. The stronger capability as general-purpose language models or other prompting methods might be important to parse semantically complex instructions into the executable sequential plan.
Compared to Section 6.1, RCI and Synapse cannot parse reverse-order instructions into plans correctly (Table 4). This can be because base-task exemplars in the prompt just have “left-to-right” instructions from base MiniWoB tasks; strongly conditioning LMAs to process instructions in a linear order while compositionality could imply a non-linear processing of instructions. The mismatch between the prompt and tasks could cause the performance drops. LMAs still fail to select correct action types (AdaPlanner) or XPath (Synapse), and they also predict unnecessary actions (RCI).
@@
Figure 4: Average success rate of large language models with advanced LLMs (gpt-4 for RCI and Synapse, and text-davinci-003 for AdaPlanner). The lighter color represents the performance in MiniWoB, the medium color does in CompWoB with gpt-3.5-turbo, and the darker color does with gpt-4 or text-davinci-003. The more capable models (gpt-4 as a generalist and text-davinci-003 as a coder) can improve the success rate of prompted LMAs but still struggle to generalize to compositional tasks (e.g. 56.0% by RCI) or to deal with reverse-order instructions (e.g. 43.5% by RCI). This may indicate that we need much better LLMs to realize deployable agents in the complex real world.